The pH of a 0,40 M formic acid, HCOOH (aq) solution is 2.08. Calculate the percentage of HCOOH molecules that are dissociated.


Select one:
a. 1,1 %
b. 11,2 %
c. 2,1 %
d. 0,89 %
1
Expert's answer
2020-08-27T05:04:49-0400

The concentration of formic acid that is in dissociated form can be determined from the pH value. As pH = -log[H+], the concentration of dissociated formic acid equals:

[H+] = 10-pH = 10-2.08 = 0,0083 M.

From here, the percentage of molecules that are dissociated equals:

w = (0.0083 / 0.4) × 100% = 2.1%


Answer: 2.1%